

********************************************************************
UARS WEEKLY STATUS REPORT - 01 OCT 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset and occultation data this past
week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE was turned on after the yaw and began successful operations on
September 28, 1999. All engineering parameters are nominal and the science
channel outputs are within expected values.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
17 days CLA (7.0) PROD (L1-L3) (473R-286C-16F=171L)
0 days GPI (4.05) PROD (CLAES L3B) (191R-164C-2F=25L)
12 days SEN (5.0) PROD (12R-12C=0L)
The CDHF has been requested by Flight Dynamics Facility (FDF) to reprocess
some 1992 data that they need for analysis and archiving of UARS Attitude
data.
DAAC DATA:
Data created and cataloged in the month of July 1999 has been previewed
for HALOE, PEM, NMC, and UKMO.
Data created for MLS Version 5 is being previewed.
UARS CDHF REHOST:
Received a "PASS" status processing Attitude Extraction Software
V7.3 in Production Test Mode for UARS DAY 1534 on CDHFB.
Level 0 Ingest from tape also successfully completed. We have reviewed
and compared the files from these two jobs with those on the current CDHF
as well as validating UCSS catalog contents. No problems have been found.
The CDHF is performing reboots of REHOST system CDHFA/CDHFB as scheduled on
Wednesdays to check system settings, logicals, etc. at startup as changes
continue to be made on these systems.
We will be starting PI Software testing this month. The HALOE new delivery,
then SUSIM are our intial candidates. The CDHF needs to recompile and relink
all PI Software prior to testing. Individual PI Software Reps will be notified
of testing results so they may reveiw output products also.
CDHF pending software upgrades: none
***********
FOT
***********
REPORTING PERIOD
This report covers 17 September, 1999 (Orbit 43822, GMT Day 261) through 24
September 1999 (Orbit 43925, GMT Day 267).
SPACECRAFT OPERATION
The observatory is now in Foward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 264/2225 43880 Off
PEM 265/1425 43890 Calibration
The beta angle ranged between the angles of 27.4 degrees to 1.2 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
50 degrees on 12 October 1999 (DOY 285).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.7 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 17 September 1999 and 24 September 1999 are provided in Appendix D.
The clock error ranged from -4.9 to -9.2 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% HALOE = 01
% PEM = 01
% FOT = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/242 - 2287.496953
% Transponder B: 99/243 - 2287.499030
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/266 show the observatory data loss to be 24 hours, 55
minutes, 26.068 seconds (an increase of 8.960 seconds since the last report
period). This is a 0.0355 percent data loss which equals a 99.9645 percent
data capture for the mission.
The increase of 8.960 seconds occurred over one (2) day:
7 8.736 seconds on Day 260 due to an IP line hit.
7 0.224 seconds on Day 264 due to an IP line hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A discrepancy was discovered between the newly delivered FDF 3-day High Gain
Gimbal Angles file (HGA), and the 4-week User Antenna View files (UAV) that it
had been receiving from FDF over the past several weeks. Investigation
revealed that the UAV files recieved from FDF had been generated using a "non-
yaw" source file resulting in post-maneuver TDRS view periods in these files to
be incorrect. The impact to operations was minimal, however the FOT had to
scramble to reschedule TDRSS events and meet schedules. Some nominal contacts
were lost due to the delay in the new TDRSS requests but ample coverage was
met.
YEAR 2000 OPERABILITY SUPPORT
The FOT received a new "Y2K" 8-week SSPP Orbital Events file from FDF this
week. The previous SSPP file did not contain any "SUN" view periods from the
beginning of the file (1999, day 365) through year 2000, day 006.
Unfortunately, the purpose of this SSPP file is to allow the FOT to schedule
the various SSPP-related operations of the spacecraft via the CMS. Without the
necessary SUN view periods, the FOT could not perform these steps... thus
preventing the testing of a large percentage of the CMS processing. The FOT
plans on processing the new SSPP file as soon as the above mentioned TDRS
problem is resolved.
MEETINGS
The FOT attended a SERBS meeting at Goddard Corporate park. Among other
mission topics UARS Y2k and hardware topics were discussed.
FUTURE SIGNIFICANT EVENTS
A meeting to discuss the UARS Safehold configuration and concerns for the MPS
and its batteries is planned for Wednesday, October 6. This meeting will be a
closed meeting with only the FOT and specialized engineering support attending.
PEM requested to have the timing in their turn on sequence tightened up in yet
another area. This is expected to be accomplished within the next few weeks.
The FOT plans on working with FDF to devise a plan to verify the integrity of
the UAV files in an effort to prevent the above mentioned situation from
happening in the future.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 September 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43880 264/2225 99-325 HALOE Power OFF
43890 265/1425 99-326 PEM PEM CALIBRATION
43920 267/1352 99-327 MACS YAW MANUEVER
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 September 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI HRDI PSE QUIET BUS CURR .851 43891
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 September 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 September 1999
17 Sept 1999 - GMT Day 260 Beta = 27.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ----- ------- ------
1 -112.0/-134.4 -2.73/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +33.6/-22.4 +6.44/+5.15 +30.4/-15.2 24.5 1.022 22.0
3 +67.2/-16.8 +5.15/+3.87 +29.6/-16.4 24.5 1.025 21.8
24 Sept 1999 - GMT Day 267 Beta = 1.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- ---------- ---- ------ ------
1 -112.0/-134.4 -3.05/-4.62 +0.0/+0.0 16.5 0.000 0.0
2 +33.6/-22.4 +5.47/+4.51 +29.6/-14.4 24.3 1.014 21.6
3 +33.6/-33.6 +4.51/+3.23 +28.8/-15.2 24.3 1.016 21.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 September 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-058 FOT 43827 261/0856 14170 Generic Late Acquisition
99-059 FOT 43827 261/0856 N/A TAC Crashed
99-058 FOT 43833 261/1906 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-055 FOT 43768 257/1029 N/A TAC Crashed
99-056 FOT 43778 258/0336 N/A TAC 8 Crashed
99-057 FOT 43809 260/0439 21348 PTP FAILURE, REBOOTED
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-058 FOT see list Sept 99 14170 Generic Late Acq.
SEVNTFW99/39